# LingoQA: Video Question Answering for Autonomous Driving

### v1

@@ Abstract
Autonomous driving has long faced a challenge with public acceptance due to the lack of explainability in the decision-making process. Video question-answering (QA) in natural language provides the opportunity for bridging this gap. Nonetheless, evaluating the performance of Video QA models has proved particularly tough due to the absence of comprehensive benchmarks. To fill this gap, we introduce LingoQA [^1] https://github.com/wayveai/LingoQA, a benchmark specifically for autonomous driving Video QA. The LingoQA trainable metric demonstrates a 0.95 Spearman correlation coefficient with human evaluations. We introduce a Video QA dataset of central London consisting of 419k samples that we release with the paper. We establish a baseline vision-language model and run extensive ablation studies to understand its performance.
@@ 1 Introduction
Communication plays a pivotal role in naturally fostering trust among individuals. However, establishing trust between users and agents remains a significant challenge within the field of artificial intelligence. Recent studies have indicated that articulating explicit reasoning steps can significantly enhance user confidence [1], in addition to improving the capabilities of Large Language Models (LLMs) [52]. The need for explainability remains critical, particularly in safety-critical domains where technology adoption hinges upon this factor [31].
@@
Consider the domain of end-to-end autonomous driving [11], where the driving policy is often executed through deep neural networks processing camera inputs to generate control commands. Despite their efficiency, these systems have grappled with the persistent challenge of providing transparent and interpretable decisions. Integrating Vision-Language Models (VLMs) into the field of autonomous driving holds the promise of enhancing user trust in these systems. Recent strides in VLMs have solidified transformers as multimodal learners, showcasing remarkable performance in tasks such as Visual Question Answering (VQA) and underscoring their proficiency in acquiring robust representations for complex tasks [14].
@@
Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust. We introduce LingoQA, a benchmark designed for autonomous driving video QA, utilizing a novel dataset comprising more than 419k QA pairs. Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving video QA, encompassing reasoning and action justifications. Additionally, we publish a comprehensive evaluation suite consisting of 1,000 examples. At the core of our benchmark lies a novel evaluation metric based on a learned text classifier called Lingo-Judge, inspired by GPT-Judge used in TruthfulQA [35]. We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [42], CIDEr [49], and GPT-4 [41] on our benchmark, while being fast enough for frequent runs during training and development. The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking of autonomous driving explainability.
@@
Equipped with this evaluation toolkit, we conducted a comprehensive empirical study on key components and their ablations in VLMs for autonomous driving. Our findings in Section 5 indicate that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13], on both Action and Scenery datasets. This process involves using 5 video frames over 4 seconds and a late video fusion technique. Our collective work, spanning the LingoQA benchmark, the visual instruction-tuning dataset, and the innovative evaluation metric, aims to propel the domain of explainable autonomous driving, laying a robust foundation for subsequent research and development endeavors. To summarise the main contributions of this paper:
@@
  • LingoQA Baseline: Through testing of various video-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique. We establish a new baseline for this field with an identified model combination. Example outputs from the model are shown in Figure LABEL:fig:lingo_examples.
@@ 2.1 Language in Autonomous Driving
Modern autonomous vehicle software relies heavily on artificial intelligence models [6, 20, 21, 23]. This, together with the increased number of such vehicles on the road, poses a fundamental challenge in terms of interpretability in the decision-making process [4]. Understanding why a decision is made is crucial for understanding areas of uncertainty, building trust, enabling effective human-AI collaboration, and ensuring safety [54]. In a survey conducted by Partners for Automated Vehicle Education (PAVE) in 2020 [1], 60% of participants stated that they would trust AVs more if they better understood the underlying process of the system. To establish trust with the general public, the inner workings of systems must be explained in a human-interpretable way, such as through language and visual explanations.
The field of autonomous driving has been embracing the opportunity to make end-to-end driving models more explainable using language, understanding that methods based purely on attention [31] are not sufficient [27]. The early explorations of GPT3.5 [46, 39] and GPT4-V [53] on autonomous driving scenarios show that LLMs/VLMs demonstrate superior performance in scene understanding and causal reasoning compared to existing autonomous systems. Works such as ADAPT [29] and LLM-Driver [10] propose multi-task learning frameworks for jointly predicting language and control outputs. Inspired by progress in large language models [47, 13, 60, 41], vision-language models [50, 7, 51, 57, 58, 59, 12, 3, 32, 41, 36, 16] and multi-modal transformers for robotics [19, 9, 8]. Closely related to our proposed baseline is DriveGPT [56], proposing a multi-modal vision-language-action model that tokenizes videos, as well as text and control actions.
@@ 2.2 Evaluation Metrics
Progress has been relatively slow for developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [30, 29, 56]. A key challenge consists of automated, reproducible evaluation metrics that are highly correlated with human ratings, particularly due to the inherent complexities in assessing natural language. ADAPT [29] reports human feedback in addition to standard natural language generation metrics, while DriveGPT [56] reports ChatGPT ratings. Automated methods such as BLEU [42], METEOR [5], ROUGE [34] show weak alignment with human feedback [49]. CIDEr [49] is also based on n-gram level similarity, as opposed to capturing the correctness of an answer based on its meaning. Newer evaluation metrics using ChatGPT have shown improvement in the area of sentence understanding, while still having limitations, such as providing high scores to elaborate, eloquent, but incorrect sentences [2]. Evaluation based on human feedback is subjective and difficult to reproduce. In this work, we address this challenge by introducing a novel video QA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
@@ 2.3 Datasets for Autonomous Driving
Recent advances in generative AI have been underpinned by training with increasingly large and diverse internet-scale datasets. Nonetheless, when evaluated zero-shot on autonomous driving datasets, many pre-trained models such as Flamingo [3] and BLIP-2 [32] fall short of expectations. Consequently, there remains a need for high-quality training datasets for autonomous driving, as well as for reliable benchmarks. Autonomous driving datasets have been focused on commentary, as opposed to question-answer pairs [31, 55]. However, we note that action, justification is a relatively easy task for foundation models, that are capable of efficiently compressing large amounts of data [17]. The advantage of VQA is that related, but slightly different questions can be asked to truly probe the validity of the model representations. Further datasets focus on ranking the importance of elements in the scene and explaining the reasoning behind the choice [45, 38]. The datasets available for VQA have been previously constructed around existing object detection datasets [43, 18]. Datasets such as NuScenesQA [43], while useful for testing perception, contain simple language outputs of on average one word per question that do not tackle the more challenging reasoning problem.
Our proposed dataset LingoQA addresses the existing gap in autonomous driving as it contains a diverse set of questions related to driving behaviours and scenery in addition to perception questions related to object presence and positioning. This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning. Examples of the complex questions and answers existent in the dataset are provided in Appendix A. Moreover, a significant advantage of LingoQA is that the intentions and reasoning behind driving decisions are directly labeled by the drivers themselves. This provides an accurate understanding of driving dynamics.
@@ 3.2 Evaluation Metric
Inspired by the G-Eval metric [37], we used GPT-4 to evaluate answers on a larger scale. Given a question and answer pair from the test set and a model’s answer, we ask GPT-4 to evaluate whether the model’s answer corresponds to a human’s answer. Notice that it doesn’t make use of any visual input. We experimented with prompts and methods achieving good quality of judgements. We achieved the highest accuracy by employing chain-of-thought prompting where we ask GPT-4 to first come up with an evaluation strategy before grading a model’s answer. However, as shown in Table 1, this leads to increased inference time. Further details are provided in Appendix C. Unfortunately, we found GPT-4 based evaluation impractical to use as a main development and training metric due to the time required to evaluate answers on our relatively small evaluation dataset (from 13min up to 50min for a single evaluation due to the API rate limit).
@@
Given these limitations and inspired by TruthfulQA GPT-Judge [35], we pursued an alternative approach using a learned text classifier, dubbed Lingo-Judge, which estimates the correctness of model answers. We measure the correctness of model predictions as an accuracy using a small transformer-based text classifier that takes in a question, the human’s, and the model’s answer and outputs a probability that the model’s answer is correct. Please note, Lingo-Judge does not receive video input and must rely only on the supporting human’s answers. For every question, we run Lingo-Judge on all combinations of (ground-truth answer, predicted answer) and take the maximum correctness estimate, as shown in Equation 1, where $S$ is the score per sample. We found this recipe yield the best predictive power provided enough diversity of human answers in our evaluation dataset.
$S=\max_{j\in\{0,1\}}F_{Judge}(pred,ground\_truth[j])$ (1)
@@
The Spearman rank correlation coefficient of our automated metric, Lingo-Judge, with human scores is 0.95, and the Pearson correlation coefficient is 0.993. These values are notably higher compared to other natural language evaluation metrics and GPT-4, as detailed in Table 1. Our analysis demonstrates that Lingo-Judge accurately mirrors human judgments, outperforming existing metrics such as BLEU, METEOR, and CIDEr, as well as GPT-4 with and without chain-of-thoughts prompting. This indicates that Lingo-Judge can effectively serve as a proxy for human labelling, which is particularly significant given the stagnant nature of metrics in autonomous driving since the introduction of the CIDEr metric in 2015. Notably, despite their limitations, prominent models like ADAPT [29] and DriveGPT [56] still use BLEU, METEOR, and CIDEr metrics and report ChatGPT ratings without analyzing their correlation to human preferences. Our work fills this gap by providing a reliable benchmark that better reflects human preferences.
@@ 3.3 Datasets
We created a collection of datasets for explainable driving. The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz. The total size of the dataset is about 10x larger than BDD-X [31], as shown in Table 3. Compared to prior datasets such as NuScenesQA [43], our dataset contains reasoning pairs in addition to object presence, description, and localisation. The answers are also more free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA. Examples of question answers pairs from LingoQA are shown in Appendix A.
@@
The scenery dataset was built to complement the action dataset by focusing on perception-related question in addition to driving behaviours. The dataset was made by densely and thoroughly labelling three 30-minute driving sessions with the ELAN video annotation software[40]. For the entire duration of the driving sessions, we provided short captions in about 15 different categories:
@@
  • Observations about relevant static road elements such as traffic lights, traffic islands, lane and intersection structures
@@ 4 Model Methodology
We propose LingoQA Baseline, a vision language model for autonomous driving based on Vicuna v1.5 [13] with 7B parameters that takes explainability beyond simple image captioning and can answer reasoning questions grounded in video outputs. We train a model that consumes a short video segment and produces answers to autonomous driving-related questions.
@@ 4.1 Architecture
The LingoQA Baseline model architecture is based on recent VLMs [32, 36, 19] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as is shown in Figure 3.
@@ 4.2 Training Recipe
Our training uses a two-step approach to better utilise video features and improve learning in answering questions based on video data. Through this two-step training, we aim for a better understanding and use of video data in addressing challenges in autonomous driving. During both stages, we train only the attention layers of both CLIP and the large language model, as well as the parameters of the Q-Former and language projection layer, while keeping all other parameters frozen. Further details regarding training parameters are presented in Appendix E.
@@ 5 Empirical Evaluation on LingoQA
With the highly modular architecture of VLMs, the question remains what architectural components of the LingoQA Baseline model and dataset composition contribute the most to its performance? We conduct several ablation studies around the architecture and training paradigm described in Section 4. To this end, we investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models. The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers. Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix F.
@@ 5.1 Training Strategy
The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance. It becomes apparent that fine-tuning is essential to yield answers relevant to autonomous driving. The model fine-tuned on the LingoQA dataset has nearly double the performance of the dataset that is pre-trained on generic VQA datasets. This shows that while generalised pre-training leads to improvements, task-specific fine-tuning is still required for optimal performance.

### v2

@@ Abstract
We introduce LingoQA, a novel dataset and benchmark for video question answering in autonomous driving. The dataset contains 28K unique scenarios and 419K annotations. Evaluating state-of-the-art vision-language models on our benchmark shows that their performance is below human capabilities, with GPT-4V responding truthfully to 56.67% of the questions compared to 93.4% for humans. For evaluation, in addition to conducting a human study, we propose a truthfulness classifier, called Lingo-Judge, that achieves a 0.95 Spearman correlation coefficient to human evaluations, surpassing existing techniques like METEOR, BLEU, CIDEr, and GPT-4. We establish a baseline vision-language model and run extensive ablation studies to understand its performance. We release our dataset and benchmark [^1] https://github.com/wayveai/LingoQA and hope that it will provide a thorough evaluation platform for future vision-language models in autonomous driving.
{strip}
Figure 1: LingoQA is a comprehensive benchmark for Video Question Answering in autonomous driving. Our baseline vision-language model on this benchmark, can answer questions related to driving reasoning, object recognition, action justification, and scene description.
@@ 1 Introduction
Communication plays a pivotal role in naturally fostering trust among individuals. However, establishing trust between users and agents remains a significant challenge within the field of artificial intelligence. Recent studies have indicated that articulating explicit reasoning steps can significantly enhance user confidence [1], in addition to improving the capabilities of Large Language Models (LLMs) [52]. The need for textual justifications remains critical, particularly in safety-critical domains where technology adoption hinges upon this factor [29].
@@
Consider the domain of end-to-end autonomous driving [11], where the driving policy is often executed through deep neural networks processing camera inputs to generate control commands. Recent strides in VLMs have solidified transformers as multimodal learners, showcasing remarkable performance in tasks such as Visual Question Answering (VQA) and underscoring their proficiency in acquiring robust representations for complex tasks [14]. Integrating Vision-Language Models (VLMs) into the field of autonomous driving holds the promise of enhancing user trust in these systems.
@@
Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust. We introduce LingoQA, a benchmark designed for autonomous driving video QA, utilizing a novel dataset comprising more than 419k QA pairs. Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving video QA, encompassing reasoning and action justifications. Additionally, we publish a comprehensive evaluation suite consisting of 1,000 examples. At the core of our benchmark lies a novel evaluation metric based on a learned text classifier called Lingo-Judge, inspired by GPT-Judge used in TruthfulQA [34]. We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [40], CIDEr [49], and GPT-4 [39] on our benchmark, while being fast enough for frequent runs during training and development. The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking video question answering in autonomous driving.
@@
Equipped with this evaluation toolkit, we conducted a comprehensive empirical study on key components and their ablations in VLMs for autonomous driving. Our findings in Section 5 indicate that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13], on both Action and Scenery datasets. This process involves using 5 video frames over 4 seconds and a late video fusion technique. Our collective work, spanning the LingoQA benchmark, the visual instruction-tuning dataset, and the innovative evaluation metric, aims to propel the domain of language-augmented autonomous driving, laying a robust foundation for subsequent research and development endeavors. To summarise the main contributions of this paper:
@@
  • LingoQA Baseline: Through testing of various video-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique. We establish a new baseline for this field with an identified model combination. Example outputs from the model are shown in Figure LingoQA: Video Question Answering for Autonomous Driving.
@@ 2.1 Language in Autonomous Driving
Modern autonomous vehicle software relies heavily on artificial intelligence models [6, 18, 19, 21]. This, together with the increased number of such vehicles on the road, poses a fundamental challenge in terms of interpretability in the decision-making process [4]. Understanding why a decision is made is crucial for understanding areas of uncertainty, building trust, enabling effective human-AI collaboration, and ensuring safety [54]. In a survey conducted by Partners for Automated Vehicle Education (PAVE) in 2020 [1], 60% of participants stated that they would trust AVs more if they better understood the underlying process of the system. To establish trust with the general public, the systems must be explained in a human-interpretable way, such as through language and visual explanations [4].
The field of autonomous driving has been embracing the opportunity to make driving models more trustworthy for their users using visual attention methods [25] or textual explanations [29]. The early explorations of GPT3.5 [45, 37] and GPT4-V [53] on autonomous driving scenarios show that LLMs/VLMs demonstrate superior performance in scene understanding and causal reasoning compared to existing autonomous systems. Works such as ADAPT [27] and LLM-Driver [10] propose multi-task learning frameworks for jointly predicting language and control outputs. Inspired by progress in large language models [47, 13, 60, 39], vision-language models [50, 7, 51, 57, 58, 59, 12, 3, 30, 39, 35, 15] and multi-modal transformers for robotics [17, 9, 8] our work incorporates language to autonomous driving. Closely related to our proposed baseline is DriveGPT [56], proposing a multi-modal vision-language-action model that tokenizes videos, as well as text and control actions.
@@ 2.2 Evaluation Metrics
Progress has been relatively slow in developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [28, 27, 56]. A key challenge consists of automated, reproducible evaluation metrics that are highly correlated with human ratings, particularly due to the inherent complexities in assessing natural language. ADAPT [27] reports human feedback in addition to standard natural language generation metrics, while DriveGPT [56] reports ChatGPT ratings. Automated methods such as BLEU [40], METEOR [5], ROUGE [33] show weak alignment with human feedback [49]. CIDEr [49] is also based on n-gram level similarity, as opposed to capturing the correctness of an answer based on its meaning. Newer evaluation metrics using ChatGPT have shown improvement in the area of sentence understanding, while still having limitations, such as providing high scores to elaborate, eloquent, but incorrect sentences [2]. Evaluation based on human feedback is subjective and difficult to reproduce. In this work, we address this challenge by introducing a novel video QA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
@@ 2.3 Datasets for Autonomous Driving
Recent advances in generative AI have been underpinned by training with increasingly large and diverse internet-scale datasets. [3] [30] This has brought into light the need for evaluation benchmarks and datasets that focus not only on specific tasks, but on reasoning areas, such as descriptive and predictive reasoning. [41] Prior works, such as the CausalVQA benchmark [32] and the Perception Test [41], a comprehensive benchmark for vision-language foundation models, probe the validity of the model representations through question answering. Autonomous driving datasets have been focused on commentary [29, 55] or constructed around existing object detection datasets [42, 16]. Datasets such as NuScenesQA [42], contain simple language outputs of on average one word per question that do not tackle the more challenging reasoning problem.
Our proposed dataset LingoQA addresses the existing gap in autonomous driving as it contains a diverse set of questions related to driving behaviours and scenery in addition to perception questions related to object presence and positioning. The evaluation dataset probes areas such as description, counting, localisation, anticipation, attention, and action justification. This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning. Examples of the complex questions and answers existent in the dataset are provided in Appendix A.
@@ 3.2 Evaluation Metric
Inspired by the G-Eval metric [36], we used GPT-4 to evaluate answers on a larger scale. Given a question and answer pair from the test set and a model’s answer, we ask GPT-4 to evaluate whether the model’s answer corresponds to a human’s answer. Notice that it does not make use of any visual input. We experimented with prompts and methods achieving good quality of judgements. We achieved the highest accuracy by employing chain-of-thought prompting where we ask GPT-4 to first come up with an evaluation strategy before grading a model’s answer. However, as shown in Table 1, this leads to increased inference time. Further details are provided in Appendix C. Unfortunately, we found GPT-4 based evaluation impractical to use as a main development and training metric due to the time required to evaluate answers on our relatively small evaluation dataset (from 13min up to 50min for a single evaluation due to the API rate limit).
@@
Given these limitations and inspired by TruthfulQA GPT-Judge [34], we pursued an alternative approach using a learned text classifier, dubbed Lingo-Judge, which estimates the correctness of model answers. We measure the correctness of model predictions as an accuracy using a small transformer-based text classifier that takes in a question, the human’s, and the model’s answer and outputs a probability that the model’s answer is correct. Please note, Lingo-Judge does not receive video input and must rely only on the supporting human’s answers. For every question, we run Lingo-Judge on all combinations of (ground-truth answer, predicted answer) and take the maximum correctness estimate, as shown in Equation 1, where $S$ is the score per sample. We found that this recipe yields the best predictive power provided enough diversity of human answers in our evaluation dataset.
$S=\max_{j\in\{0,1\}}F_{\mathrm{Judge}}(\mathrm{prediction},\mathrm{ground\_truth}[j])$ (1)
@@
The Spearman rank correlation coefficient of our automated metric, Lingo-Judge, with human scores is 0.95, and the Pearson correlation coefficient is 0.993. These values are considerably higher compared to other natural language evaluation metrics and GPT-4, as detailed in Table 1. Our analysis demonstrates that Lingo-Judge accurately mirrors human judgments, outperforming existing metrics such as BLEU, METEOR, and CIDEr, as well as GPT-4 with and without chain-of-thought prompting. This indicates that Lingo-Judge can effectively serve as a proxy for human labelling, which is particularly significant given the stagnant nature of metrics in autonomous driving since the introduction of the CIDEr metric in 2015. Notably, despite their limitations, prominent models like ADAPT [27] and DriveGPT [56] still use BLEU, METEOR, and CIDEr metrics and report ChatGPT ratings without analyzing their correlation to human preferences. Our work fills this gap by providing a reliable benchmark that better reflects human preferences.
@@ 3.3 Datasets
We created a collection of datasets for bringing language to autonomous driving. The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz. The total size of the dataset is about 10x larger than BDD-X [29], as shown in Table 3. Compared to prior datasets such as NuScenesQA [42], our dataset contains reasoning pairs in addition to object presence, description, and localisation. The answers are also more free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA. Examples of question answers pairs from LingoQA are shown in Appendix A.
@@
The scenery dataset was built to complement the action dataset by focusing on perception-related questions in addition to driving behaviours. The dataset was made by densely and thoroughly labelling three 30-minute driving sessions with the ELAN video annotation software[38]. For the entire duration of the driving sessions, we provided short captions in about 15 different categories:
@@
  • Observations about relevant road elements such as traffic lights, traffic islands, lane and intersection structures
@@ 4 Model Methodology
We propose LingoQA Baseline, a vision language model for autonomous driving based on Vicuna v1.5 [13] with 7B parameters that can answer reasoning questions grounded in video outputs. We train a model that consumes a short video segment and produces answers to autonomous driving-related questions.
@@ 4.1 Architecture
The LingoQA Baseline model architecture is based on recent VLMs [30, 35, 17] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as shown in Figure 3.
@@ 4.2 Training Recipe
Our training uses a two-step approach to better utilise video features and improve learning when answering questions based on video data. Through this two-step training, we aim to better understand and use video data. In the first stage, we train the self-attention layers for the LLM and the vision encoder (QKV), together with all the Q-Former parameters and the linear language projection layer. In the second stage, we fine-tune the same parameters as in the previous stage, keeping the vision encoder frozen. Further details regarding training parameters are presented in Appendix F.
@@ 5 Empirical Evaluation on LingoQA
With the highly modular architecture of VLMs, the question remains what architectural components of the LingoQA Baseline model and dataset composition contribute the most to its performance? We conduct several ablation studies around the architecture and training paradigm described in Section 4. To this end, we investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models. The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers. Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix G.
@@ 5.1 Training Strategy
The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance. It becomes apparent that fine-tuning leads to improved answers relevant to autonomous driving. The model fine-tuned on the LingoQA dataset has double the performance of the dataset that is pre-trained on generic VQA datasets.